OpenStudy (anonymous):

Explain why (–4x)^0 = 1, but –4x^0 = –4

OpenStudy (anonymous):

-4x^0 = -4 times x^0 = -4 times 1 = -4 anything to the 0 (except 0 and infinity) is 1. Calculus deals with the 0 and infinity cases. so (-4x)^0 is like a^0 = 1, where a = -4x

OpenStudy (anonymous):

I am confused.

OpenStudy (anonymous):

How do you know that a=-4x?

OpenStudy (anonymous):

(-4x)^3 = -64x^3 (-4x)^2 = 16x^2 (-4x)^1 = -4x^1 if we increase the exponent by 1 we multiply by -4x. likewise, if we decrease the exponent by 1 we divide by -4x. what is -4x divided by -4x?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

yep. does that one makes sense?

OpenStudy (anonymous):

Oh I think I get it now, if it was 4x^0 then it would be 1, right?

OpenStudy (anonymous):

not quite... when written like that, x is the only thing being raised to the exponent. 4x^2 = 4*x*x, only the x is exponentiated. do you remember the order of operations? (PEMDAS)? exponents come first them multiplication. and that's what is going on there. 4 is a multiplier and x is the base. with (-4x)^0 the base is -4x, the whole thing.

OpenStudy (anonymous):

What do you mean the whole thing?

OpenStudy (anonymous):

i mean -4x is the base. so (-4x)^2 = (-4x)*(-4x) = 16x^2 when you have -4x^2 = -4*x^2=-4*x*x

OpenStudy (anonymous):

i mean the -4x is the base, not just the x. that's what i meant by the whole thing. hopefully that's a little clearer.

OpenStudy (anonymous):

hmm, I think I get it know... Thanks, you earned yourself a medal and fan :)

OpenStudy (anonymous):

thanks for your patience!!!

OpenStudy (jdoe0001):

\(\large {\bf ({\color{brown}{ -4x}})^0\to {\color{brown}{ 1}} \\ \quad \\ -4{\color{brown}{ x}}^0\to -4\cdot {\color{brown}{ x}}^0\to -4\cdot {\color{brown}{ 1}}\to -4 }\)